please help me to solve this problem
The bermuda triangle supposedly causes trouble for aircraft pilots . it has a perimeter of 3075 mi . The shortest side measures 75 mi less than the middle side ,and the longest side measures 375 mi more than the middle side .Find the lengths of the three sides
my email shiela.balan@yahoo.com

Answer by nerdybill(7384)

The bermuda triangle supposedly causes trouble for aircraft pilots . it has a perimeter of 3075 mi . The shortest side measures 75 mi less than the middle side ,and the longest side measures 375 mi more than the middle side .Find the lengths of the three sides
.
Let m = measure of middle side
then from "shortest side measures 75 mi less than the middle side"
m-75 = measure of shortest side
and from "longest side measures 375 mi more than the middle side"
m+375 = measure of longest side
.
sum of the sides = perimeter
m + m-75 + m+375 = 3075
3m -75 +375 = 3075
3m + 300 = 3075
m + 100 = 1025
m = 925 miles (middle side)
.
shortest side:
m-75 = 925-75 = 850 miles
.
longest side:
m+375 = 75+375 = 450 miles